# SceneCOT: Eliciting Grounded Chain-of-Thought Reasoning in 3D Scenes

### v2

@@ Abstract
Existing research on 3D Large Language Models (LLMs) still struggles to achieve grounded question-answering, primarily due to the under-exploration of the mechanism of human-like scene-object grounded reasoning. This paper bridges the gap by presenting a novel framework. We first introduce a grounded Chain-of-Thought reasoning method in 3D scenes (SceneCOT), decoupling a complex reasoning task into simpler and manageable problems, and building corresponding visual clues based on multimodal expert modules. To enable such a method, we develop SceneCOT-185K, the first large-scale grounded CoT reasoning dataset, consisting of 185K high-quality instances. Extensive experiments across various complex 3D scene reasoning benchmarks demonstrate that our new framework achieves strong performance with high grounding-QA coherence. To the best of our knowledge, this is the first successful application of CoT reasoning to 3D scene understanding, enabling step-by-step human-like reasoning and showing potential for extension to broader 3D scene understanding scenarios.
@@ 1 Introduction
Understanding 3D scenes is a fundamental capability for building human-level embodied agents (Chen et al., 2019; Yang et al., 2025b; Chen et al., 2024a; Song et al., 2025). Despite growing interest and progress in this area (Chen et al., 2024c; Jia et al., 2024; Zhu et al., 2024b; Huang et al., 2024; Chen et al., 2023b; Zhou et al., 2024), reasoning in complex 3D environments remains highly challenging. Unlike image-based reasoning (Guo et al., 2023; Ou et al., 2024; Chen et al., 2024d; Liu et al., 2023), 3D reasoning—especially in embodied and situated settings (Ma et al., 2023; Linghu et al., 2024)—requires navigating large spaces, interpreting intricate spatial relations, and coping with partial observability. Addressing these challenges calls for principled task decomposition and step-by-step reasoning, yet existing research has largely overlooked this aspect.
Recent benchmarks highlight the consequences of this gap. In particular, Beacon3D (Huang et al., 2025) reveals that current 3D vision-language models often produce plausible answers without grounding them in the scene, leading to poor grounding–QA coherence. This underlines a fundamental limitation: while models may generate fluent responses, they fail to connect intermediate grounding steps with the final reasoning outcome. We argue that achieving human-like 3D reasoning requires answers that emerge from transparent, grounded, and stepwise reasoning processes.
@@
In this work, we present SceneCOT, a novel framework for step-by-step grounded reasoning in 3D scenes. SceneCOT explicitly decomposes complex 3D reasoning tasks into four stages: (1) task recognition and analysis, (2) task-relevant region localization, (3) entity and attribute grounding with multimodal expert modules, and (4) grounded reasoning that integrates intermediate results into coherent final answers. This hierarchical workflow ensures that each answer is supported by explicit grounding steps, thereby enhancing grounding–QA coherence. To enable such reasoning, we construct SceneCOT-185K, the first large-scale grounded CoT dataset for 3D reasoning, containing over 185K high-quality reasoning traces across diverse 3D QA benchmarks. Each trace captures the full stepwise reasoning trajectory, including task-oriented region selection, object grounding, and final answer generation. Extensive experiments on situated reasoning benchmark MSQA (Linghu et al., 2024), as well as grounding–QA coherence evaluation on Beacon3D (Huang et al., 2025), show that SceneCOT achieves strong performance while producing reasoning that is both interpretable and faithfully grounded.
@@
We demonstrate that SceneCOT achieves strong performance on challenging 3D reasoning benchmarks, and importantly, improves grounding–QA coherence, as revealed by in-depth analyses on Beacon3D.
@@ 2 Related Work
Understanding 3D scenes is essential for developing human-like intelligence, and recent studies have increasingly explored the use of LLMs for 3D vision-language understanding. Early efforts such as 3D-LLM (Hong et al., 2023) lift multi-view 2D features into 3D space and align them with text embeddings, while PointLLM (Xu et al., 2023) leverages point cloud encoders for object-level geometry reasoning. LEO (Huang et al., 2024) aligns object-centric 3D representations with LLMs for 3D vision-language (3D-VL) tasks. Subsequent works have expanded this line of research (Chen et al., 2024b; Yang et al., 2025a; Fu et al., 2025; Chu et al., 2024; Zhang et al., 2024d; Qi et al., 2023). For instance, Grounded 3D-LLM (Chen et al., 2024c) enhances point-level semantic alignment with language, Chat-Scene (Zhang et al., 2024a) builds per-object 3D features via mask proposals, LLaVA-3D (Zhu et al., 2024a) integrates 3D positional embeddings into LLaVA (Liu et al., 2023), Video-3D LLM (Zheng et al., 2024) leverages pretrained video MLLMs for temporal-spatial context, and SplatTalk (Thai et al., 2025) applies 3D Gaussian Splatting (Kerbl et al., 2023) to generate language-aligned 3D tokens. Despite these advances, most 3D-LLMs rely on sparse supervision in an end-to-end training paradigm, with limited exploration of intermediate reasoning processes. As shown in Beacon3D (Huang et al., 2025), these models often produce answers that appear correct but lack explicit links to scene grounding, leading to poor grounding–QA coherence. This reveals a fundamental gap: while 3D-LLMs can encode multimodal representations, they rarely incorporate structured, step-by-step reasoning mechanisms that are essential for robust 3D scene understanding. We argue that Chain-of-Thought (CoT) reasoning remains largely untapped in 3D, but is necessary for overcoming overfitting and building more interpretable, human-like 3D-VL models.
@@
LLMs have demonstrated remarkable reasoning abilities across diverse domains such as mathematics, programming, and scientific QA (Shao et al., 2024b; Jaech et al., 2024). These capabilities are significantly enhanced by Chain-of-Thought (CoT) prompting (Wei et al., 2022), which decomposes complex tasks into step-by-step subproblems. Inspired by this, researchers have sought to extend reasoning into multimodal settings. Flamingo (Alayrac et al., 2022) bridges frozen vision and language backbones with cross-attention, Shikra (Chen et al., 2023a) and KOSMOS-2 (Peng et al., 2023) emphasize grounded reasoning over visual inputs, while OMG-LLaVA (Zhang et al., 2024b) unifies pixel-, object-, and region-level understanding. More recent works attempt to bring step-by-step reasoning into MLLMs. V (Wu & Xie, 2024) incorporates sequential visual search for fine-grained recognition, Video-of-Thought (Fei et al., 2024) introduces a perception-to-recognition workflow for video reasoning, and Visual CoT (Shao et al., 2024a) explicitly integrates bounding boxes and patch-level grounding as intermediate “thoughts.” Commercial models such as GPT-o3 and GPT-o3 mini (OpenAI, 2025) showcase impressive visual reasoning abilities by combining multiple visual experts. However, these advances remain confined to 2D inputs. They lack explicit grounding in 3D space, limiting their ability to reason about spatially complex, embodied scenarios. In summary, while CoT reasoning has transformed both text-only and 2D multimodal models, it has not yet been fully explored in 3D scene understanding. Bridging this gap requires designing frameworks that integrate step-by-step grounded reasoning with explicit grounding–QA coherence, which is the focus of our work.
@@ 4.1 Metadata Collection
First, we design a rule-based procedure to extract target objects based on the agent’s location and orientation within the 3D scene. This begins by parsing the question and extracting directional cues using regular expression matching. In MSQA, directional information typically falls into two categories: cardinal directions (left, right, front, back) and clock-based directions (e.g., “at the 1–12 o’clock”). For instance, given the question “How many tables are on my right?”, we extract all objects located to the right of the agent. This rule-based method ensures that answers can be accurately inferred from the object list within the corresponding sub-region. Secondly, for the target object entities, we design a rule-based method for Existence/Counting to ensure the correctness. For the remaining sub-tasks, we inherit the official annotations of the target objects in the released data.
@@ 5.1 Experimental Setup
Tasks and Data. We evaluate SceneCOT on two representative 3D scene reasoning tasks: Situated Reasoning (Situated Reasoning) and Object-Centric Reasoning (Object-Centric Reasoning). For Situated Reasoning, we adopt the MSQA benchmark built on ScanNet, following the evaluation protocol of (Thai et al., 2025). To ensure fair and consistent comparison, we use the refined Version-2.1 of MSQA (Linghu et al., 2024), and re-implement MSR3D and GPT-4o on this version. For MSR3D, we follow the official setup, which includes the merged training data and object mask proposals from Mask3D. For Object-Centric Reasoning, we evaluate on the Beacon3D benchmark (Huang et al., 2025), which explicitly measures grounding–QA coherence. Following the official settings, we use ground-truth object masks for all methods to isolate reasoning capability from detection performance.
Evaluation Protocols and Baselines. On both MSQA and Beacon3D, we use GPT-score as the primary evaluation metric. For MSQA, models jointly predict object masks and textual answers, while for Beacon3D, evaluation is performed using ground-truth masks. Beacon3D further reports two complementary metrics: (1) GPT-Score (by case), which averages scores across individual QA pairs, and (2) GPT-Score (by object), which requires all QA pairs associated with an object to be correct, directly reflecting grounding–QA coherence. In addition, we analyze the grounding–QA coherence metrics provided in Beacon3D to highlight the improvements brought by our method. We compare SceneCOT against a broad set of 3D scene reasoning baselines. On Beacon3D, we focus on object-centric baselines that explicitly incorporate grounding, including GPT-4o (OpenAI, 2023), MSR3D (Linghu et al., 2024), PQ3D (Zhu et al., 2024b), Chat-Scene (Zhang et al., 2024a), and SceneVerse (Jia et al., 2024). On MSQA, we additionally include SplatTalk (Thai et al., 2025), which reports results on this benchmark. To ensure fairness, we retrain several baselines on our training set and report the updated results.
@@
5.2 Main Results on 3D QA and Grounding–QA Coherence
@@
In the main results, we evaluate both 3D QA performance and grounding–QA coherence, and compare SceneCOT with baseline methods.
SceneCOT demonstrates strong QA performance on Situated Reasoning and Object-Centric Reasoning. It achieves notable gains on the challenging Counting sub-task by explicitly enumerating objects in the relevant sub-region based on semantic similarity, providing clear visual grounding for reasoning. SceneCOT also performs well on Spatial tasks, including Refer and Spatial Relationship. Its performance is lower on Existence and Attribute compared to LEO and MSR3D, mainly due to grounding errors (see Sec.˜5.3). While slightly below Chat-Scene overall, SceneCOT substantially outperforms all baselines on Beacon3D thanks to its scene–object grounded reasoning. In contrast, Chat-Scene directly predicts answers from question and object-centric tokens without explicit grounding, limiting its ability on complex tasks like Counting, where object-level reasoning is critical. Additional baseline comparisons in Sec.˜C.1 further confirm the competitiveness of SceneCOT.
SceneCOT achieves strong grounding–QA coherence. On Beacon3D, SceneCOT achieves the highest Good Coherence (34.7), substantially ahead of all baselines. This demonstrates that our framework most reliably combines correct grounding with correct QA, rather than succeeding on one dimension while failing on the other. By contrast, LEO reaches a QA (Obj.) score of 7.8 but suffers from an extremely low GC of 1.6, highlighting a large mismatch between answering correctly and grounding correctly. Similarly, methods like SceneVerse (GC: 20.4, QA (Obj.): 6.6) and Chat-Scene (GC: 19.5, QA (Obj.): 7.8) show moderate QA performance but struggle to align it with consistent grounding. In addition to its advantage in GC, SceneCOT also delivers the highest QA (Obj.) score (23.2), a large margin over all other methods. This reflects not only stronger per-object QA ability but also a tighter alignment between answers and the objects they refer to. Together, these results confirm that explicitly enforcing grounding before reasoning enables SceneCOT to achieve both accurate QA and high grounding–QA coherence, setting it apart from prior baselines.
@@ 5.3 Ablation Study and Analyses
Question type recognition is essential. As illustrated in Fig.˜2, SceneCOT first identifies the question type and builds the reasoning chain accordingly. For example, the symbolic engine outputs polar coordinates for Navigation and 3D bounding boxes for Spatial Relationship. To assess its role, we conduct an ablation where all questions are forced into the same type. As shown in Fig.˜4, treating all questions “as existence” restricts the model to object probabilities only, leading to clear performance degradation. Likewise, forcing all questions to Spatial Relationship significantly reduces performance on Navigation, where polar coordinates are more suitable. These results demonstrate that recognizing the correct question type is crucial for constructing grounded reasoning chains.
Region recognition is crucial, particularly for object-centric reasoning. To enhance grounding in Situated Reasoning, SceneCOT explicitly filters out irrelevant objects, narrowing the scope of candidates to task-relevant regions. Removing this component and providing all scene objects results in a substantial performance drop, especially on Counting, Refer, and Attribute (see Fig.˜4). This confirms that region recognition not only reduces noise but also strengthens grounding–QA coherence by aligning reasoning with localized evidence.
Grounding loss plays a key role. In SceneCOT, the grounding module is trained with an additional loss term based on PQ3D, encouraging accurate object grounding during learning. Ablating this grounding loss (see Eq.˜1) yields a noticeable performance decline on MSQA, particularly for Counting, Refer, and Navigation (see Fig.˜4). Interestingly, performance on Existence is less affected, likely because semantic labels alone suffice for this type of question. Overall, these results highlight the grounding loss as an important factor for maintaining step-by-step grounding consistency and improving reasoning accuracy.
@@
Exploring the QA upper bound with better grounding information. In MSQA, predictions are limited by noisy semantic labels, masks, and object probabilities. To measure their impact, we test oracle settings that remove: Semantic errors (SE): inaccurate labels and masks; Grounding errors (GE): incorrect object probabilities. Removing SE improves performance, especially on Counting. Removing both SE and GE pushes results close to the oracle bound, with Counting/Existence near 100 and Navigation/Refer reaching 87.2/84.9. These gains confirm that accurate grounding and semantics directly translate into better answers, demonstrating strong grounding–QA coherence in our framework. Spatial Relationship remains the hardest task, as it requires mapping coordinates to natural descriptions. Overall, the analysis shows that improving label, mask, and grounding quality is key to advancing 3D reasoning.

### v3

@@ Abstract
Existing research on 3D Large Language Models (LLMs) still struggles to achieve grounded question-answering, primarily due to the under-exploration of the mechanism of human-like scene-object grounded reasoning. This paper bridges the gap by presenting a novel framework. We first introduce a grounded Chain-of-Thought reasoning method in 3D scenes (SceneCOT), decoupling a complex reasoning task into simpler and manageable problems, and building corresponding visual clues based on multimodal expert modules. To enable such a method, we develop SceneCOT-185K, the first large-scale grounded CoT reasoning dataset, consisting of 185K high-quality instances. Extensive experiments across various complex 3D scene reasoning benchmarks demonstrate that our new framework achieves strong performance with high grounding-QA coherence. To the best of our knowledge, this is the first successful application of CoT reasoning to 3D scene understanding, enabling step-by-step human-like reasoning and showing potential for extension to broader 3D scene understanding scenarios. Code and data are available on project page.
@@ 1 Introduction
Understanding 3D scenes is a fundamental capability for building human-level embodied agents (Chen et al., 2019; Yang et al., 2025b; Chen et al., 2024a; Gong et al., 2023; Song et al., 2025). Despite growing interest and progress in this area (Chen et al., 2024c; Jia et al., 2024; Zhu et al., 2024b; Huang et al., 2024b; Chen et al., 2023b; Zhou et al., 2024), reasoning in complex 3D environments remains highly challenging, especially in embodied and situated settings (Ma et al., 2023; Linghu et al., 2024). Reasoning in 3D scenes requires navigating large spaces, interpreting intricate spatial relations, and coping with partial observability. Addressing these challenges calls for principled task decomposition and step-by-step reasoning, yet existing research has largely overlooked this aspect.
Recent benchmarks highlight the consequences of this gap. In particular, Beacon3D (Huang et al., 2025a) reveals that current 3D vision-language models often produce plausible answers without grounding them in the scene, leading to poor grounding-QA coherence. This underlines a fundamental limitation: while models may generate fluent responses, they fail to connect intermediate grounding steps with the final reasoning outcome. We argue that achieving human-like 3D reasoning requires answers that emerge from transparent, grounded, and stepwise reasoning processes.
@@
In this work, we present SceneCOT, a novel framework for step-by-step grounded reasoning in 3D scenes. SceneCOT explicitly decomposes complex 3D reasoning tasks into four stages: (1) task recognition and analysis, (2) task-relevant region localization, (3) entity and attribute grounding with multimodal expert modules, and (4) grounded reasoning that integrates intermediate results into coherent final answers. This hierarchical workflow ensures that each answer is supported by explicit grounding steps, thereby enhancing grounding-QA coherence. To enable such reasoning, we construct SceneCOT-185K, the first large-scale grounded CoT dataset for 3D reasoning, containing over 185K high-quality reasoning traces across diverse 3D QA benchmarks. Each trace captures the full stepwise reasoning trajectory, including task-oriented region selection, object grounding, and final answer generation. Extensive experiments on situated reasoning benchmark MSQA (Linghu et al., 2024), as well as grounding-QA coherence evaluation on Beacon3D (Huang et al., 2025a), show that SceneCOT achieves strong performance while producing reasoning that is both interpretable and faithfully grounded.
@@
We demonstrate that SceneCOT achieves strong performance on challenging 3D reasoning benchmarks, and importantly, improves grounding-QA coherence, as revealed by in-depth analyses on Beacon3D.
@@ 2 Related Work
Understanding 3D scenes is essential for developing human-like intelligence, and recent studies have increasingly explored the use of LLMs for 3D vision-language understanding. Early efforts such as 3D-LLM (Hong et al., 2023) lift multi-view 2D features into 3D space and align them with text embeddings, while PointLLM (Xu et al., 2023) leverages point cloud encoders for object-level geometry reasoning. LEO (Huang et al., 2024b) aligns object-centric 3D representations with LLMs for 3D vision-language (3D-VL) tasks. Subsequent works have expanded this line of research (Chen et al., 2024b; Yang et al., 2025a; Fu et al., 2025; Chu et al., 2024; Zhang et al., 2024c; Qi et al., 2023). For instance, Grounded 3D-LLM (Chen et al., 2024c) enhances point-level semantic alignment with language, Chat-Scene (Huang et al., 2024a) builds per-object 3D features via mask proposals, LLaVA-3D (Zhu et al., 2024a) integrates 3D positional embeddings into LLaVA (Liu et al., 2023), Video-3D LLM (Zheng et al., 2024) leverages pretrained video MLLMs for temporal-spatial context, and SplatTalk (Thai et al., 2025) applies 3D Gaussian Splatting (Kerbl et al., 2023) to generate language-aligned 3D tokens. LEO-VL (Huang et al., 2025b) further improves the efficiency of scene representation for 3D MLLMs. Despite these advances, most 3D-LLMs rely on sparse supervision in an end-to-end training paradigm, with limited exploration of intermediate reasoning processes. As shown in Beacon3D (Huang et al., 2025a), these models often produce answers that appear correct but lack explicit links to scene grounding, leading to poor grounding-QA coherence. This reveals a fundamental gap: while 3D-LLMs can encode multimodal representations, they rarely incorporate structured, step-by-step reasoning mechanisms that are essential for robust 3D scene understanding. We argue that Chain-of-Thought (CoT) reasoning remains largely untapped in 3D, but is necessary for overcoming overfitting and building more interpretable, human-like 3D-VL models.
@@
LLMs have demonstrated remarkable reasoning abilities across diverse domains such as mathematics, programming, and scientific QA (Guo et al., 2023; Ou et al., 2024; Chen et al., 2024d; Huang et al., 2022; Shao et al., 2024b; Jaech et al., 2024). These capabilities are significantly enhanced by Chain-of-Thought (CoT) prompting (Wei et al., 2022), which decomposes complex tasks into step-by-step subproblems. Inspired by this, researchers have sought to extend reasoning into multimodal settings. Flamingo (Alayrac et al., 2022) bridges frozen vision and language backbones with cross-attention, Shikra (Chen et al., 2023a) and KOSMOS-2 (Peng et al., 2023) emphasize grounded reasoning over visual inputs, while OMG-LLaVA (Zhang et al., 2024a) unifies pixel-, object-, and region-level understanding. More recent works attempt to bring step-by-step reasoning into MLLMs. V (Wu and Xie, 2024) incorporates sequential visual search for fine-grained recognition, Video-of-Thought (Fei et al., 2024) introduces a perception-to-recognition workflow for video reasoning, and Visual CoT (Shao et al., 2024a) explicitly integrates bounding boxes and patch-level grounding as intermediate “thoughts.” Commercial models such as GPT-o3 and GPT-o3 mini (OpenAI, 2025) showcase impressive visual reasoning abilities by combining multiple visual experts. However, these advances remain confined to 2D inputs. They lack explicit grounding in 3D space, limiting their ability to reason about spatially complex, embodied scenarios. In summary, while CoT reasoning has transformed both text-only and 2D multimodal models, it has not yet been fully explored in 3D scene understanding. Bridging this gap requires designing frameworks that integrate step-by-step grounded reasoning with explicit grounding-QA coherence, which is the focus of our work.
@@ 4.1 Metadata Collection
First, we design a rule-based procedure to extract target objects based on the agent’s location and orientation within the 3D scene. This begins by parsing the question and extracting directional cues using regular expression matching. In MSQA, directional information typically falls into two categories: cardinal directions (left, right, front, back) and clock-based directions (e.g., “at the 1-12 o’clock”). For instance, given the question “How many tables are on my right?”, we extract all objects located to the right of the agent. This rule-based method ensures that answers can be accurately inferred from the object list within the corresponding sub-region. Secondly, for the target object entities, we design a rule-based method for Existence/Counting to ensure the correctness. For the remaining sub-tasks, we inherit the official annotations of the target objects in the released data.
@@ 5.1 Experimental Setup
Tasks and Data. We evaluate SceneCOT on two representative 3D scene reasoning tasks: Situated Reasoning (Situated Reasoning) and Object-Centric Reasoning (Object-Centric Reasoning). For Situated Reasoning, we adopt the MSQA benchmark built on ScanNet, following the evaluation protocol of (Thai et al., 2025). To ensure fair and consistent comparison, we use the refined Version-2.1 of MSQA (Linghu et al., 2024), and re-implement MSR3D and GPT-4o on this version. For MSR3D, we follow the official setup, which includes the merged training data and object mask proposals from Mask3D. For Object-Centric Reasoning, we evaluate on the Beacon3D benchmark (Huang et al., 2025a), which explicitly measures grounding-QA coherence. Following the official settings, we use ground-truth object masks for all methods to isolate reasoning capability from detection performance.
Evaluation Protocols and Baselines. On both MSQA and Beacon3D, we use GPT-score as the primary evaluation metric. For MSQA, models jointly predict object masks and textual answers, while for Beacon3D, evaluation is performed using ground-truth masks. Beacon3D further reports two complementary metrics: (1) GPT-Score (by case), which averages scores across individual QA pairs, and (2) GPT-Score (by object), which requires all QA pairs associated with an object to be correct, directly reflecting grounding-QA coherence. In addition, we analyze the grounding-QA coherence metrics provided in Beacon3D to highlight the improvements brought by our method. We compare SceneCOT against a broad set of 3D scene reasoning baselines. On Beacon3D, we focus on object-centric baselines that explicitly incorporate grounding, including GPT-4o (OpenAI, 2023), MSR3D (Linghu et al., 2024), PQ3D (Zhu et al., 2024b), Chat-Scene (Huang et al., 2024a), and SceneVerse (Jia et al., 2024). On MSQA, we additionally include SplatTalk (Thai et al., 2025), which reports results on this benchmark. To ensure fairness, we retrain several baselines on our training set and report the updated results.
@@
5.2 Main Results on 3D QA and Grounding-QA Coherence
@@
In the main results, we evaluate both 3D QA performance and grounding-QA coherence, and compare SceneCOT with baseline methods.
SceneCOT demonstrates strong QA performance on Situated Reasoning and Object-Centric Reasoning. It achieves notable gains on the challenging Counting sub-task by explicitly enumerating objects in the relevant sub-region based on semantic similarity, providing clear visual grounding for reasoning. SceneCOT also performs well on Spatial tasks, including Refer and Spatial Relationship. Its performance is lower on Existence and Attribute compared to LEO and MSR3D, mainly due to grounding errors (see Sec.˜5.3). While slightly below Chat-Scene overall, SceneCOT substantially outperforms all baselines on Beacon3D thanks to its scene-object grounded reasoning. In contrast, Chat-Scene directly predicts answers from question and object-centric tokens without explicit grounding, limiting its ability on complex tasks like Counting, where object-level reasoning is critical. Additional baseline comparisons in Sec.˜C.1 further confirm the competitiveness of SceneCOT.
SceneCOT achieves strong grounding-QA coherence. On Beacon3D, SceneCOT achieves the highest Good Coherence (34.7), substantially ahead of all baselines. This demonstrates that our framework most reliably combines correct grounding with correct QA, rather than succeeding on one dimension while failing on the other. By contrast, LEO reaches a QA (Obj.) score of 7.8 but suffers from an extremely low GC of 1.6, highlighting a large mismatch between answering correctly and grounding correctly. Similarly, methods like SceneVerse (GC: 20.4, QA (Obj.): 6.6) and Chat-Scene (GC: 19.5, QA (Obj.): 7.8) show moderate QA performance but struggle to align it with consistent grounding. In addition to its advantage in GC, SceneCOT also delivers the highest QA (Obj.) score (23.2), a large margin over all other methods. This reflects not only stronger per-object QA ability but also a tighter alignment between answers and the objects they refer to. Together, these results confirm that explicitly enforcing grounding before reasoning enables SceneCOT to achieve both accurate QA and high grounding-QA coherence, setting it apart from prior baselines.
@@ 5.3 Ablation Study and Analyses
Question type recognition is essential. As illustrated in Fig.˜2, SceneCOT first identifies the question type and builds the reasoning chain accordingly. For example, the symbolic engine outputs polar coordinates for Navigation and 3D bounding boxes for Spatial Relationship. To assess its role, we conduct an ablation where all questions are forced into the same type. As shown in Fig.˜5, treating all questions “as existence” restricts the model to object probabilities only, leading to clear performance degradation. Likewise, forcing all questions to Spatial Relationship significantly reduces performance on Navigation, where polar coordinates are more suitable. These results demonstrate that recognizing the correct question type is crucial for constructing grounded reasoning chains.
Region recognition is crucial, particularly for object-centric reasoning. To enhance grounding in Situated Reasoning, SceneCOT explicitly filters out irrelevant objects, narrowing the scope of candidates to task-relevant regions. Removing this component and providing all scene objects results in a substantial performance drop, especially on Counting, Refer, and Attribute (see Fig.˜5). This confirms that region recognition not only reduces noise but also strengthens grounding-QA coherence by aligning reasoning with localized evidence.
Grounding loss plays a key role. In SceneCOT, the grounding module is trained with an additional loss term based on PQ3D, encouraging accurate object grounding during learning. Ablating this grounding loss (see Eq.˜1) yields a noticeable performance decline on MSQA, particularly for Counting, Refer, and Navigation (see Fig.˜5). Interestingly, performance on Existence is less affected, likely because semantic labels alone suffice for this type of question. Overall, these results highlight the grounding loss as an important factor for maintaining step-by-step grounding consistency and improving reasoning accuracy.
@@
Exploring the QA upper bound with better grounding information. In MSQA, predictions are limited by noisy semantic labels, masks, and object probabilities. To measure their impact, we test oracle settings that remove: Semantic errors (SE): inaccurate labels and masks; Grounding errors (GE): incorrect object probabilities. Removing SE improves performance, especially on Counting. Removing both SE and GE pushes results close to the oracle bound, with Counting/Existence near 100 and Navigation/Refer reaching 87.2/84.9. These gains confirm that accurate grounding and semantics directly translate into better answers, demonstrating strong grounding-QA coherence in our framework. Spatial Relationship remains the hardest task, as it requires mapping coordinates to natural descriptions. Overall, the analysis shows that improving label, mask, and grounding quality is key to advancing 3D reasoning.
5.4 Additional Evaluation of Grounding Performance
To more thoroughly evaluate the grounded aspect of our 3D-CoTs reasoning, we consider both traditional grounding benchmarks and QA-driven grounding benchmarks. As shown in Tab.˜4, we incorporate multiple benchmarks to comprehensively evaluate performance in both grounding-only settings (Nr3D, Beacon3D) and QA-driven grounding tasks (MSQA, SQA3D, ScanQA). For the latter, the evaluation measures the model’s ability to localize the specific object instances required to answer a question (e.g., grounding only the specific chairs to the agent’s right). We provide the details of the benchmarks in Sec.˜C.3.
| Method | Nr3D Top-1 | B3D Top-1 | MSQA F1@50 | SQA3D F1@50 | ScanQA F1@50 |
| Chat-Scene | 39 | 62.7 | 15.9 | 25.9 | 35.9 |
| PQ3D | 66 | 76 | 10.6 | 10.8 | 19.2 |
| SceneCOT | 57.7 | 67.8 | 52.1 | 51.6 | 40.8 |
The results show that SceneCOT achieves strong and balanced grounding performance across all benchmarks. It performs competitively on Nr3D and Beacon3D, and it significantly surpasses prior methods on MSQA, SQA3D, and ScanQA grounding. Notably, our model is not fine-tuned on SQA3D or ScanQA; thus, the strong zero-shot grounding performance highlights the effectiveness and generalizability of our approach. These high-quality grounding outputs provide clear and reliable visual cues that support the downstream reasoning chains.
In contrast, while baseline models like PQ3D demonstrate strong results on traditional benchmarks like Nr3D, they perform poorly on QA-driven grounding tasks. This reveals a limitation in their ability to follow complex instructions in reasoning scenarios. Chat-Scene shows relatively better grounding on SQA3D and ScanQA but still lags behind SceneCOT across the full spectrum of tasks. Overall, this evaluation demonstrates that our framework is built upon robust grounding capabilities for both spatial localization and reasoning-oriented tasks.